# Elliptic flow of deuterons from simulations with hybrid model

### v1

@@ I Introduction
Production of deuterons and larger nuclear clusters from the hot matter produced in relativistic heavy-ion collisions is not fully understood, presently. The observation that their abundances are rather well described by the grand-canonical statistical model with chemical freeze-out temperature around 150 MeV Andronic:2017pug ; Andronic:2016nof is quite puzzling, as their binding energy is smaller by two orders of magnitude. This poses a question, how far one can push an interpretation with such a model in the description of other experimental observables than the multiplicities.
@@ II Motivation
In Vozabova:2024jum , two of us have investigated the discriminating power of differential elliptic flow of clusters for the acceptance or rejection of one of the two competing production models .
Deuterons are femtoscopic probe. This is implicitly included in the formalism of coalescence and implemented since many years Mrowczynski:1987oid ; Scheibl:1998tk ; Sun:2018mqq ; Mrowczynski:1989jd ; Mrowczynski:1989jd ; Mrowczynski:1992gc ; Mrowczynski:1993cx ; Blum:2019suo ; Kachelriess:2019taq ; Kachelriess:2020amp ; Kachelriess:2023jis ; Kachelriess:2022khq In coalescence, the yield of clusters with certain momentum strongly depends on the size and density profile of the region that produces its constituent nucleons. The relevant scale to which the size is compared is the width of the cluster wave function. The yield in thermal production scales with the volume of the producing homogeneity region. The dependences on the sizes of the homogeneity regions are different in the two mechanisms.
@@ III The model
Our simulations use the chain TRENTo+vHLLE+Hadron Sampler+SMASH. We studied Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV, for centralities up to 80%.
Initial distributions of energy density and entropy density for the fluid dynamical simulations were generated by three-dimensional version of TRENTo Ke:2016jrd . Parameters were chosen Ke:2019jbh so that the best agreement for $p_{t}$ spectra and $v_{2}(p_{t})$ is achieved for protons, pions, and kaons at all centralities, as measured by ALICE collaboration ALICE:2013mez . The values of parameters are summarised in Table 1.
@@
$\sqrt{s_{NN}}$ 2.76 TeV
p 0.0
k 2.0
w 0.59
n 114.0
$\mu_{0}$ 0.0
$\sigma_{0}$ 2.9
$\gamma_{0}$ 7.3
skewness 1.0
$\mathrm{J}$ 0.75
@@
$\tau_{0}$ 0.6 fm/$c$
$\varepsilon_{\mathrm{crit}}$ 0.32 GeV/fm3
$\eta/s$ 0.12
$\zeta/s$ 0.4
@@
$\varepsilon_{\mathrm{crit}}$ 0.32 GeV/fm3
$t_{\mathrm{max}}$ 100 fm/$c$
The initial distributions were further evolved with vHLLE Karpenko:2013wva . The model assumed shear viscosity with $\eta/T=0.12$. Bulk viscosity with a maximum around critical temperature and a peak value of $\zeta/T=0.4$ was parametrised according to eq.(43) of Schenke:2020mbo , as it is also implemented in vHLLE. The fluid-dynamical simulation evolved the bulk matter down to energy density $\varepsilon_{\mathrm{crit}}=0.32$ GeV/fm3. This marked the particlisation hypersurface, at which Hadron Sampler from the vHLLE+SMASH package was used.
Each hydrodynamic evolution served as initial condition to configurations of hadrons that were further evolved with SMASH SMASH:2016zqf , version 2.2. The maximum time of the simulations was set to 100 fm/$c$.
@@ V Results
The results seem to indicate that the elliptic flow of deuterons that are thermally produced and evolved in SMASH is higher than that from coalescence. Surprisingly, this is opposite to the observation made in Vozabova:2024jum , where coalescence deuterons had higher $v_{2}$. We also note an observation which we made in our simulations, that the elliptic flow of direct thermal deuterons grows until very late times. We have noticed such a growth when comparing simulations with $t_{\mathrm{max}}=50$ fm/$c$ and $t_{\mathrm{max}}=100$ fm/$c$.
When compared to data, direct thermal production overshoots the measured $v_{2}$, while coalescence which was tuned on proton and pion production reproduces the data well.
@@ VI Conclusions
Our simulations with hybrid model including vHLLE hydrodynamics and SMASH as hadron afterburner shows qualitatively different results for the differential elliptic flow than the previous analysis of two of us Vozabova:2024jum . Hence, we conclude that $v_{2}(p_{t})$ is qualitatively not robust enough to be the decisive probe of deuteron production mechanism.
@@
Secondly, larger clusters are heavier and therefore they are suppressed in production. For example, 3He is suppressed with respect do deuterons by a factor of 300 in Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV ALICE:2015wav . In turn, this means that the number of simulated events would have to be increased at least by this factor. At the moment, this is prohibitively demanding for us.

### v2

@@ I Introduction
Production of deuterons and larger nuclear clusters from the hot matter produced in relativistic heavy-ion collisions is not fully understood, presently. The observation that their abundances are rather well described by the grand-canonical statistical model with chemical freeze-out temperature around 150 MeV Andronic:2017pug ; Andronic:2016nof is quite puzzling, as their binding energy is smaller by two orders of magnitude. Moreover, the typical size of a nucleon cluster is larger than the interparticle distance, hence the interaction of the two nucleons is expected to be screened by surrounding hadrons. Note that the hot fireball is usually expected not to decompose immediately after the chemical freeze-out but cools down further down during a few following fm/$c$. The clusters would have to exist and survive in such an environment. This poses a question, how far one can push an interpretation with such a model in the description of experimental observables other than the multiplicities.
@@ II Motivation
In Vozabova:2024jum , two of us have investigated the discriminating power of differential elliptic flow of clusters for the acceptance or rejection of one of the two alternative production models.
Deuterons are femtoscopic probe. This is implicitly included in the formalism of coalescence and implemented since many years Mrowczynski:1987oid ; Scheibl:1998tk ; Sun:2018mqq ; Mrowczynski:1989jd ; Mrowczynski:1989jd ; Mrowczynski:1992gc ; Mrowczynski:1993cx ; Blum:2019suo ; Kachelriess:2019taq ; Kachelriess:2020amp ; Kachelriess:2023jis ; Kachelriess:2022khq In coalescence, the yield of clusters with certain momentum strongly depends on the size and density profile of the region that produces its constituent nucleons. The relevant scale to which the size is compared is the width of the cluster wave function. The yield in thermal production just scales with the volume of the producing homogeneity region. The dependences on the sizes of the homogeneity regions are different in the two mechanisms.
@@ III The model
Our simulations use the chain TRENTo+vHLLE+Hadron Sampler+SMASH. We studied Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV, for centralities up to 70%.
Initial distributions of energy density and entropy density for the fluid dynamical simulations were generated by three-dimensional version of TRENTo Ke:2016jrd . Parameters were chosen Ke:2019jbh so that the best agreement for $p_{t}$ spectra and $v_{2}(p_{t})$ is achieved for protons, pions, and kaons at all centralities, as measured by ALICE collaboration ALICE:2013mez . Relative skewness ansatz for the rapidity distribution was assumed Ke:2019jbh . The values of parameters are summarised in Table 1.
@@
$\sqrt{s_{NN}}$ collision energy 2.76 TeV
$p$ generalized mean parameter 0.0
$k$ multiplicity fluctuation shape 2.0
$w$ Gaussian nucleon width 0.59 fm
$n$ normalization 114.0
$\mu_{0}$ mean rapidity shift 0.0
$\sigma_{0}$ rapidity distribution width 2.9
$\gamma_{\mathrm{rel}}$ rapidity distribution skewness 7.3
$J$ pseudorapidity Jacobian 0.75
@@
$\tau_{0}$ initial time 0.6 fm/$c$
$\varepsilon_{\mathrm{crit}}$ critical energy density 0.32 GeV/fm3
$\eta/s$ shear viscosity 0.12
$\zeta/s$ bulk viscosity 0.4
@@
$\varepsilon_{\mathrm{crit}}$ critical energy density 0.32 GeV/fm3
$t_{\mathrm{max}}$ maximum simulation time 100 fm/$c$
The initial distributions were further evolved with vHLLE Karpenko:2013wva . The model assumed shear viscosity with $\eta/s=0.12$. Bulk viscosity with a maximum around critical temperature and a peak value of $\zeta/s=0.4$ was parametrised according to eq.(43) of Schenke:2020mbo , as it is also implemented in vHLLE. The fluid-dynamical simulation evolved the bulk matter down to energy density $\varepsilon_{\mathrm{crit}}=0.32$ GeV/fm3. This marked the particlisation hypersurface, at which Hadron Sampler from the vHLLE+SMASH package was used.
Each hydrodynamic evolution served as initial condition to configurations of hadrons that were further evolved with SMASH SMASH:2016zqf , version 2.2. The maximum time in the simulations was set to 100 fm/$c$ after which any simulation is stopped.
@@
It turned out that it is important to evolve the transport simulation for sufficiently long time. Cross-section of deuteron-induced reactions, as implemented in SMASH Oliinychenko:2018ugs , are of the order of tens of mb, with $\sigma(\sqrt{s})$ for $\pi d$ scattering peaking above 200 mb. Hence, they scatter even at quite low density, as the collision criterion would still be fulfilled for distances between pion and deuteron more than 20 fm. We have examined the reliability of our results by running for different values of $t_{\mathrm{max}}$, which is the parameter in SMASH that sets the maximum running time after which the evolution is stopped in any case. By simulating for $t_{\mathrm{m}ax}=50$ fm/$c$, 100 fm/$c$, and 150 fm/$c$ we checked that our results change between 50 and 100 fm/$c$, but stay unchanged between 100 and 150 fm/$c$. Further on we report the results obtained for 100 fm/$c$.
@@ V Results
The slight underprediction of spectra normalization actually agrees qualitatively with the observation made in ALICE:2015wav by the ALICE collaboration. When looking at the deuteron-to-proton ratio as a function of the number of participating nucleons (Fig. 16 in ALICE:2015wav ), it slightly grows when going from central to non-central collisions. Because of the size of the error bars, which are also due to extrapolation of the $p_{t}$ spectra to full $p_{t}$ coverage, the dependence is also consistent with a constant. Here, in Fig. 4 the error bars are much smaller and our theoretical curve lies clearly below the experimental data points.
@@
The results seem to indicate that the elliptic flow of deuterons that are thermally produced and evolved in SMASH is higher than that from coalescence. This feature appears rather systematically in all simulated centrality classes. Surprisingly, this is opposite to the observation made in Vozabova:2024jum , where coalescence deuterons had higher $v_{2}$.
When compared to data, direct thermal production overshoots the measured $v_{2}$, while coalescence which was tuned on proton and pion production reproduces the data well. Nevertheless, coalescence slightly overshoots the data in the more central classes of events. A closer inspection of proton elliptic flow, plotted in Fig. 3, shows that it is also the case for protons (and even pions) with transverse momenta below 1 GeV/$c$. Thus, since the deuteron consists of two nucleons it is reasonable to expect that the tension with data reaches up to twice that upper limit.
@@ VI Conclusions
Our simulations with hybrid model including vHLLE hydrodynamics and SMASH as hadron afterburner show qualitatively different results for the differential elliptic flow than the previous analysis of two of us Vozabova:2024jum : elliptic flow here is larger from the thermal production than from coalescence, while it was the other way around in Vozabova:2024jum . We are thus prevented from stating that deuteron elliptic flow is always larger in one mechanism in comparison to the other. Differential elliptic flow cannot be the decisive probe to distinguish the two mechanisms in such a simple way.
It is the difference between what is called thermal model here and what it was in Vozabova:2024jum that is the cause for the different results. Let us compare the two models more closely. To this end, recall that in Vozabova:2024jum all hadrons were produced from an extended blast-wave parametrization. This included resonances which were let to decay. No rescattering between hadrons happened after that.
In both cases of the thermal model, the assumption is that the deuteron is a pointlike particle. This is an important feature because in thermal models with expansion the unique assignment of position leads to the determination of the momentum through the connection between position and the flow velocity. (Note that coalescence takes the finite size of the deuteron implicitly into account.) The difference between Vozabova:2024jum and here is that there the clusters are immediately produced in the final state while here they are generated at the beginning of the transport simulation. After that they still rescatter and can be destroyed as well as created. Note that they have large cross-section for the interaction with pions and the large abundance of the latter ones. As we already mentioned, the rescattering goes on until rather late time and contributed to the buildup of the elliptic flow. The higher $v_{2}$ is a natural consequence of this. From this point of view, the two thermal models, as implemented here and in Vozabova:2024jum are rather different. Nevertheless, what we use here is the closest implementation of such a model that can be consistently run within the framework of a hybrid model.
@@
Secondly, larger clusters are heavier and therefore they are suppressed in production. For example, 3He is suppressed with respect to deuterons by a factor of 300 in Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ TeV ALICE:2015wav . In turn, this means that the number of simulated events would have to be increased at least by this factor. At the moment, this is prohibitively demanding for us.